OpenStudy (anonymous):

Suppose that a NASCAR race car is moving to the right with a constant velocity of +82 m/s. What is the average acceleration of the car? (b) Twelve seconds later, the car is halfway around the track and traveling in the opposite direction with the same speed. What is the average acceleration of the car? I understand (a) but I don't understand how to solve for (b) can someone please help me?

OpenStudy (anonymous):

average acceleration = average velocity/time = 82 / 12 = 41 / 6 = 6. 83

OpenStudy (anonymous):

is my answer correct

OpenStudy (anonymous):

Acceleration = dv/dt thus it is the change in velocity over the change in time or the derivative of the velocity time graph. Thus if you have a constant velocity (including both speed AND direction) then you have 0 acceleration.

OpenStudy (anonymous):

in part b you have to realize that velocity includes direction. Velocity is a vector. Because of this, when the car is going in the opposite direction, its velocity is now -82 m/s

OpenStudy (anonymous):

so, in those 12 seconds the car went from moving at +82 m/s to -82 m/s. The difference which is equivalent to the change is velocity is obviously 82*2=164 m/s

OpenStudy (anonymous):

so to find average acceleration we look change in velocity over change in time = 164m/s/12s = 13.67 m/s^2